# FineRadScore: A Radiology Report Line-by-Line Evaluation Technique Generating Corrections with Severity Scores

### v1

@@ Abstract
The current gold standard for evaluating generated chest x-ray (CXR) reports is through radiologist annotations. However, this process can be extremely time-consuming and costly, especially when evaluating large numbers of reports. In this work, we present FineRadScore, a Large Language Model (LLM)-based automated evaluation metric for generated CXR reports. Given a candidate report and a ground-truth report, FineRadScore gives the minimum number of line-by-line corrections required to go from the candidate to the ground-truth report. Additionally, FineRadScore provides an error severity rating with each correction and generates comments explaining why the correction was needed. We demonstrate that FineRadScore’s corrections and error severity scores align with radiologist opinions. We also show that, when used to judge the quality of the report as a whole, FineRadScore aligns with radiologists as well as current state-of-the-art automated CXR evaluation metrics. Finally, we analyze FineRadScore’s shortcomings to provide suggestions for future improvements. Code to run FineRadScore can be found here.
@@ 1 Introduction
Automating the evaluation of AI-generated reports is a challenging task. The current gold standard is manual evaluation by radiologists, yet evaluating a large number of reports in this manner is costly and time-consuming, creating a need for automated metrics. Researchers have explored natural language generation (NLG) metrics such as BLEU (Papineni et al. (2002)) and BERTScore (Zhang et al. (2019)) to compare the generated report with a ground-truth report. However, these metrics were designed for non-medical text and fail to capture important features of medical text due to the lack of domain-specific knowledge needed to evaluate radiology reports (Yu et al. (2023), Liu et al. (2023)). For example, if a model generated a sentence such as “there is a focal lesion” and the ground-truth report stated “there is no focal lesion,” this pair would score highly on BLEU-1, even though radiologists would consider it a poor match.
@@ 4.2 Annotation Process
After obtaining these 2 reports for each study, we recruited 2 radiologists to annotate subsets of the selected radiology studies. Both radiologists were given a 30-minute orientation, as well as instructions and examples to complete their annotation tasks.
@@ 5 Experiments
5. Results 1-4 hold when the candidate report is stylistically different from the ground-truth report.
@@ 6.2 FineRadScore Text Rewrite and Insertion Quality
Dataset Model BLEU-1 Scores BLEU-2 Scores
ReFiSco-v1 GPT-4 0.8601 0.8191
Opus 0.8590 0.8101
ReFiSco-v0 GPT-4 0.7776 0.7398
Opus 0.8371 0.8095
@@ 6.3 Correcting Candidate Reports Using FineRadScore
We now aim to ensure that after applying FineRadScore’s corrections to our candidate report, we obtain a corrected report that is more aligned with our ground-truth report. Therefore, we went through each correction outputted by FineRadScore and applied it to the original line in the candidate report to obtain a corrected report. We then compare the FineRadScore-corrected report and the ground-truth report in three ways. First, we compare the corrected candidate report and the ground-truth report’s text similarities using BLEU-2 scores. However, as noted in the introduction, lexical similarity need not be correlated with semantic alignment. Therefore, we use RadGraph-F1 (Jain et al. (2021)) and RadCliQ (Yu et al. (2023)), two evaluation metrics designed for chest x-ray reports and which have been shown to align well with radiologists. RadGraph-F1 measures the similarity between the sets of medically relevant words and their relationships for a pair of reports and outputs a score. The score ranges from 0 to 1, with 1 representing the highest alignment between the reports. RadCliQ is a composite metric that combines both lexical and clinical metrics for a pair of reports to output a score. The lower the score, the more aligned the pair of reports are.
@@
As shown in Table 3, when we apply the corrections generated by FineRadScore to our candidate report, we obtain higher BLEU scores when comparing this corrected report to our ground-truth report. This finding is consistent across models and datasets and is most prominent with FineRadScore-Opus. FineRadScore also produces corrected candidate reports with higher RadGraph-F1 scores and lower RadCliQ scores. Therefore, according to these metrics, FineRadScore is producing higher-quality reports that are more meidcally aligned with the ground-truth reports.
@@ 6.4 FineRadScore Clinical Severity Ratings
We next study whether FineRadScore is able to estimate the clinical severity of each error it corrects. Following Tian et al. (2023), we treat unwanted comparisons to priors as not actionable errors in quantitative analyses. When no corrections are found, we set the clinical severity rating to 0. Then, for each of the four severity ratings, we assign the numbers 1 to 4 in order of increasing severity (not actionable, actionable nonurgent error, urgent error, emergent error). We take the maximum clinical severity over all lines in a given report to produce a clinical severity score. We then compare FineRadScore’s report-level clinical severity score with the radiologist’s.
To provide a baseline for how well FineRadScore captures the severity rating for each line, we also measure how often different radiologists agree with each other, using inter-rater disagreement to provide a baseline. Inter-rater disagreement gives us a sense of the best we can expect from FineRadScore. For example, if all raters find the same error but are split between whether the error is emergent or urgent, the best we can expect from FineRadScore is to find the same error and rate it within that range of severities. In ReFiSco-v1, we did not collect annotations from different radiologists for a single report, so we cannot compute inter-rater disagreement. However, in ReFiSco-v0, for generations with multiple radiologist ratings, we choose one of the raters to be the candidate report at random and treat all other ratings as ground-truth reports. We then compare the candidate radiologist’s severity ratings to the ground truth radiologists’ to obtain inter-rater disagreement scores on ReFiSco-v0.
@@ 7.1 A Shuffled Ground Truth
Dataset Model BLEU-1 Scores BLEU-2 Scores
ReFiSco-v1 GPT-4 0.8807 0.8303
Opus 0.8223 0.7634
ReFiSco-v0 GPT-4 0.7504 0.7030
Opus 0.8074 0.7533
@@ 7.2 A Paraphrased Candidate Report
Dataset Model BLEU-1 Score BLEU-2 Scores
ReFiSco-v1 GPT-4 0.7060 0.6640
Opus 0.6224 0.5847
@@ 8 Alignment with Radiologists on ReXVal
9 Qualitative Analysis of FineRadScore
@@ 10 Limitations
Future work may focus on improving performance when the ground truth differs significantly stylistically from the candidate report. As shown in Section 9, FineRadScore is still prone to identifying corrections based off of syntactic differences instead of clinically significant differences. Even so, it is able to achieve comparable performance to state of the art evaluation techniques such as RadCliQ (Yu et al. (2023)) and G-Rad (Chaves et al. (2024)). Therefore, we recommend future work focus on finding a method that can solely identify clinically relevant errors, while ignoring differences in phrasing. We hypothesize that such a finding would improve the performance of LLM-based evaluation techniques across the board.
@@
Ethical Considerations
In all of our experiments, we used Azure’s OpenAI endpoints to access GPT-4. We used Anthropic’s endpoints to access Claude-3 Opus. We also signed and adhered to the PhysioNet Data Use Agreement to ensure the confidentiality of the MIMIC-CXR data.
@@ Appendix A Prompt Instructions
We started off each prompt with the same set of prompt instructions, which is a redacted version of the instructions we gave the radiologists. Both few-shot and zero-shot prompts began with these set of instructions. These instructions are shown below.

### v2

@@ Abstract
The current gold standard for evaluating generated chest x-ray (CXR) reports is through radiologist annotations. However, this process can be extremely time-consuming and costly, especially when evaluating large numbers of reports. In this work, we present FineRadScore, a Large Language Model (LLM)-based automated evaluation metric for generated CXR reports. Given a candidate report and a ground-truth report, FineRadScore gives the minimum number of line-by-line corrections required to go from the candidate to the ground-truth report. Additionally, FineRadScore provides an error severity rating with each correction and generates comments explaining why the correction was needed. We demonstrate that FineRadScore’s corrections and error severity scores align with radiologist opinions. We also show that, when used to judge the quality of the report as a whole, FineRadScore aligns with radiologists as well as current state-of-the-art automated CXR evaluation metrics. Finally, we analyze FineRadScore’s shortcomings to provide suggestions for future improvements.
@@ 1 Introduction
Automating the evaluation of AI-generated reports is a challenging task. The current gold standard is manual evaluation by radiologists, yet evaluating a large number of reports in this manner is costly and time-consuming, creating a need for automated metrics. Researchers have explored natural language generation (NLG) metrics such as BLEU (Papineni et al. (2002)) and BERTScore (Zhang et al. (2019)) to compare the generated report with a ground-truth report. However, these metrics were designed for non-medical text and fail to capture important features of medical text due to the lack of domain-specific knowledge needed to evaluate radiology reports (Yu et al. (2023), Liu et al. (2023), Adams et al. (2023)). For example, if a model generated a sentence such as “there is a focal lesion” and the ground-truth report stated “there is no focal lesion,” this pair would score highly on BLEU-1, even though radiologists would consider it a poor match.
@@
While it can also be used as an intermediate step to generate preference data for RLHF frameworks, FineRadScore needs both the candidate and ground truth reports as input. Therefore, it should still primarily be used as an evaluation metric.
@@ 4.2 Annotation Process
After obtaining these 2 reports for each study, we recruited 2 radiologists to annotate subsets of the selected radiology studies. One of the radiologists is board-certified with 7 years of experience, and the other is a resident. Both radiologists were given a 30-minute orientation, as well as instructions and examples to complete their annotation tasks.
@@ 5 Experiments
5. Results 1-4 hold when the candidate report is stylistically different (shuffled or paraphrased) from the ground-truth report.
@@ 6.2 FineRadScore Text Rewrite and Insertion Quality
Dataset Model BLEU-1 Scores BLEU-2 Scores BERT Scores
ReFiSco-v1 GPT-4 0.8601 0.8191 0.9503
Opus 0.8590 0.8101 0.9418
ReFiSco-v0 GPT-4 0.7776 0.7398 0.8980
Opus 0.8371 0.8095 0.9175
@@ 6.3 Correcting Candidate Reports Using FineRadScore
We now aim to ensure that after applying FineRadScore’s corrections to our candidate report, we obtain a corrected report that is more aligned with our ground-truth report. In particular, we wanted to see if a FineRadScore-corrected report exhibited 1) more text similarity and 2) more semantic similarity to the ground truth reports. Therefore, we went through each correction outputted by FineRadScore and applied it to the original line in the candidate report to obtain a corrected report. We then compare the FineRadScore-corrected report and the ground-truth report in three ways. First, we compare the corrected candidate report and the ground-truth report’s text similarities using BLEU-2 scores. However, as noted in the introduction, lexical similarity need not be correlated with semantic alignment. Therefore, we use RadGraph-F1 (Jain et al. (2021)) and RadCliQ (Yu et al. (2023)), two evaluation metrics designed for chest x-ray reports and which have been shown to align well with radiologists. RadGraph-F1 measures the similarity between the sets of medically relevant words and their relationships for a pair of reports and outputs a score. The score ranges from 0 to 1, with 1 representing the highest alignment between the reports. RadCliQ is a composite metric that combines both lexical and clinical metrics for a pair of reports to output a score. The lower the score, the more aligned the pair of reports are.
@@
As shown in Table 3, when we apply the corrections generated by FineRadScore to our candidate report, we obtain higher BLEU scores when comparing this corrected report to our ground-truth report. This finding is consistent across models and datasets and is most prominent with FineRadScore-Opus. FineRadScore also produces corrected candidate reports with higher RadGraph-F1 scores and lower RadCliQ scores. Therefore, according to these metrics, FineRadScore is producing higher-quality reports that are more medically aligned with the ground-truth reports.
@@ 6.4 FineRadScore Clinical Severity Ratings
We next study whether FineRadScore is able to estimate the clinical severity of each error it corrects. Following Tian et al. (2023), we treat unwanted comparisons to priors as not actionable errors in quantitative analyses. When no corrections are found, we set the clinical severity rating to 0. Then, for each of the four severity ratings, we assign the numbers 1 to 4 in order of increasing severity (not actionable, actionable nonurgent error, urgent error, emergent error). We note that there are multiple valid ways to get a report-level score from line-level scores and that our paper explores both taking the maximum line-level score and summing all line-level scores. In this section, we take the maximum clinical severity over all lines in a given report to produce a clinical severity score. We then compare FineRadScore’s report-level clinical severity score with the radiologist’s.
To provide a baseline for how well FineRadScore captures the severity rating for each line, we also measure how often different radiologists agree with each other, using inter-rater disagreement to provide a baseline. Inter-rater disagreement gives us a sense of the best we can expect from FineRadScore. For example, if all raters find the same error but are split between whether the error is emergent or urgent, the best we can expect from FineRadScore is to find the same error and rate it within that range of severities. In ReFiSco-v1, we did not collect annotations from different radiologists for a single report, so we cannot compute inter-rater disagreement. However, in ReFiSco-v0, for generations with multiple radiologist ratings, we randomly choose one of the radiologist’s ratings to be the candidate and treat all other radiologists’ ratings as ground-truth ratings. We then compare the candidate radiologist’s severity ratings to the ground truth radiologists’ to obtain inter-rater disagreement scores on ReFiSco-v0.
@@ 7.1 A Shuffled Ground Truth
Dataset Model BLEU-1 Scores BLEU-2 Scores BERT Scores
ReFiSco-v1 GPT-4 0.8807 0.8303 0.9473
Opus 0.8223 0.7634 0.9239
ReFiSco-v0 GPT-4 0.7504 0.7030 0.8793
Opus 0.8074 0.7533 0.9071
@@ 7.2 A Paraphrased Candidate Report
Dataset Model BLEU-1 Score BLEU-2 Scores BERT Scores
ReFiSco-v1 GPT-4 0.7060 0.6640 0.9014
Opus 0.6224 0.5847 0.8827
@@ 8 Alignment with Radiologists on ReXVal
Qualitative Analysis of FineRadScore
@@ 9 Limitations
Existing automated evaluation metrics such as RadGraph and RadCliQ are currently being used to evaluate FineRadScore. Ideally, we would have a radiologist evaluate the FineRadScore-corrected reports and ensure that they are indeed more aligned with the ground truth report and would recommend future works to do this. However, previously established metrics such as RadCliQ and RadGraph-F1 consistently indicate that the FineRadScore-corrected reports are more aligned with the ground truth. As a result, we believe that this is a strong indication that the corrections are indeed creating reports that are more aligned with the ground truth.
Future work may also focus on improving performance when the ground truth differs significantly stylistically from the candidate report. As shown in the qualitative analysis, FineRadScore is still prone to identifying corrections based off of syntactic differences instead of clinically significant differences. Even so, it is able to achieve comparable performance to state of the art evaluation techniques such as RadCliQ (Yu et al. (2023)) and G-Rad (Chaves et al. (2024)). Therefore, we recommend future work focus on finding a method that can solely identify clinically relevant errors, while ignoring differences in phrasing. We hypothesize that such a finding would improve the performance of LLM-based evaluation techniques across the board.
@@
Finally, we recommend expanding upon the ReFiSco-v1 dataset by collecting duplicate annotations, where multiple radiologists annotate the same report. In doing so, we can use inter-rater disagreement as a baseline, just as we did on ReFiSco-v0 in Section 6.4.
Code Availability
In all of our experiments, we used Azure’s OpenAI endpoints to access GPT-4. We used Anthropic’s endpoints to access Claude-3 Opus. We also signed and adhered to the PhysioNet Data Use Agreement to ensure the confidentiality of the MIMIC-CXR data. Code to run FineRadScore can be found here.
@@ 10 Conclusion
[…]
Many thanks to Shirley Shen and Kathy Yu for their help! We also acknowledge support in the form of azure credits from the Microsoft Accelerating Foundation Models Research (AFMR) grant.
@@
The candidate reports were split into lines using the following regular expression:
[…]
@@ Appendix A Prompt Instructions
We started off each prompt with the same set of prompt instructions, which is a redacted version of the instructions we gave the radiologists. These examples were selected with the goal of covering a wide variety of possible errors in the candidate report, as well as the fact that purely stylistic differences should be ignored. Both few-shot and zero-shot prompts began with these set of instructions. These instructions are shown below.